0708661:

Solve the following equation for x: 6(4x+5)=3(x+8)+3. Round to the nearest hundredth.

savavge:

i solved ur propblem

darkknight:

distribute, then solve for x

darkknight:

don't give the answer directly

darkknight:

6(4x+5)=3(x+8)+3 24x+30 = 3x +24 +3 24x+30=3x+27 21x = -3 x= -3/21 = -1/7